Find the equation of the line through (2,3) and perpendicular to 3x-2y=7?
Answer by mananth(16946)

3 x -2 y = 7
Find the slope of this line

-2 y = -3 x + 7
Divide by -2
y = 2 x + -3.5
Compare this equation with y=mx+b
slope m = 3/2

The slope of a line perpendicular to the above line will be the negative reciprocal -2/3
m1*m2=-1
The slope of the required line will be - 2/3

m= - 2/3 ,point ( 2 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = -1.33 + b
b= 4.33
m= - 2/3
The required equation is y = - 2/ 3 x + 13/ 3